Contents
- 1 What is the probability of getting a 3 on a die?
- 2 What is the probability that all three?
- 3 What is the probability of getting 3 on the dice and head on the coin when the dice is rolled and a coin is tossed simultaneously?
- 4 What is the probability of getting at most two tails?
- 5 What is the probability of getting all tails?
- 6 What’s the probability of a 3 on a dice?
- 7 How is the probability of getting a 7 calculated?
What is the probability of getting a 3 on a die?
Two (6-sided) dice roll probability table
| Roll a… | Probability |
|---|---|
| 3 | 3/36 (8.333%) |
| 4 | 6/36 (16.667%) |
| 5 | 10/36 (27.778%) |
| 6 | 15/36 (41.667%) |
What is the probability that all three?
So, assuming the dice are ‘fair’ (that each of the six numbers has a probability of 1/6 of showing up on each of the dice), there is a probability of 1/36 that all three dice will show the same number. If 3 dice are rolled, the total possible outcomes (a, b, c) = 6x6x6 = 6^(3) .
What is the probability of getting 3 on the dice and head on the coin when the dice is rolled and a coin is tossed simultaneously?
1/4
N=3: To get 3 heads, means that one gets only one tail. This tail can be either the 1st coin, the 2nd coin, the 3rd, or the 4th coin. Thus there are only 4 outcomes which have three heads. The probability is 4/16 = 1/4.
What is the probability of getting 7?
Probability of getting a number <7 is 66=1.
What is the probability of rolling a sum of 3?
We divide the total number of ways to obtain each sum by the total number of outcomes in the sample space, or 216. The results are: Probability of a sum of 3: 1/216 = 0.5% Probability of a sum of 4: 3/216 = 1.4%
What is the probability of getting at most two tails?
Hence, the probability of getting at most 2 tails is 7/8.
What is the probability of getting all tails?
If the coin is fair, then the odds of getting heads or tails should be equal, 12. Then 3 tosses of tails will have a chance of 12⋅12⋅12=18.
What’s the probability of a 3 on a dice?
Only one side of the dice is a ‘3’. So the probability of rolling a three is 1 / 3 . The number three makes up 1 out of 6 sides of the dice and on average will be rolled once every six rolls. A dice has six equally likely outcomes: 1, 2, 3, 4, 5 and 6. The probability of rolling each number is 1 out of 6.
What is the probability of getting 7 on a 10 sided die?
There is a simple relationship – p = 1/s, so the probability of getting 7 on a 10 sided die is twice that of on a 20 sided die. The probability of rolling the same value on each die – while the chance of getting a particular value on a single die is p, we only need to multiply this probability by itself as many times as the number of dice.
How to calculate the probability of rolling a die?
The probability of rolling the same value on each die – while the chance of getting a particular value on a single die is p, we only need to multiply this probability by itself as many times as the number of dice. In other words, the probability P equals p to the power n, or P = pⁿ = (1/s)ⁿ.
How is the probability of getting a 7 calculated?
The probability of getting a given value for the total on the dice may be calculated by taking the total number of ways that value can be produced and dividing it by the total number of distinguishable outcomes. So the probability of a 7 on the dice is 1/6 because it can be produced in 6 ways out of a total of 36 possible outcomes.